OpenStudy (mackenzie2013):

Write the equation that describes the line in slope-intercept form. slope = 4, point (3, −2) is on the line y = 4x + 14 y = 4x − 2 y = 4x + 10 y = 4x − 14

OpenStudy (kenrick143):

Use point-slope form.

OpenStudy (kenrick143):

(y-ysub1)= m(x-xsub1) y-(-2)=4(x-3) ---> simplify y+2=4x-12 y=4x-12-2 y=4x-14 ---> final answer. Where the point would pass through and having a slope of 4